OpenStudy (nali):

I need help plz i do not get this at all What is the kinetic energy, in J, of: (A) an Ar atom moving with a speed of 650m/s? (B) a mole of Ar atoms moving with a speed of 650m/s (Hint: 1 amu =1.66 * 10^-27 kg)?

OpenStudy (nali):

@Sheng @Preetha @oOKawaiiOo plz anyone help

OpenStudy (anonymous):

From my Chem book, Nali: \[KE = 1/2 mv ^{2}\]

OpenStudy (nali):

well i know that but i need help finding the mass like last time

OpenStudy (anonymous):

So, KE = 1/2 (the mass of one atom of Ar)(650 m/s)^2

OpenStudy (anonymous):

Did you find the mass of a single atom of Ar from the last question?

OpenStudy (nali):

yes i found it it is 6.63 x 10^ -26 kg

OpenStudy (anonymous):

KE=1/2 (it is 6.63 x 10^ -26 kg)(650 m/s)^2 I'll do my math...

OpenStudy (nali):

i need help with part 2

OpenStudy (anonymous):

Isn't one mole of Ar 39.948 g? So it would follow KE = 1/2 (39.948 g)(650 m/s)^2. Please double check my work, though.

OpenStudy (nali):

no u have to use both amu and avargado's number

OpenStudy (anonymous):

I'm sorry that wasn't correct. I'll be following this so that I can get the correct answer, too.

OpenStudy (nali):

it is ok thanku for trying